For each of the following pairs of compounds, describe a simple chemical or physical test to distinguish between them, stating specific observations.

  • CH3CH2CH2OH and CH3CHOHCH3


  • CH3CH2CH=CH2 AND CH3CH=CHCH3
1
Expert's answer
2021-03-11T08:40:55-0500

1)CH3-CHOH-CH3 + HCl —ZnCl2—> CH3-CHCl-CH3↓ + H2O

CH3-CH2-CH2OH doesn't react.

2) CH2=CH2-CH2-CH3 + KMnO4 + H2SO4 → CH3-CH2-COOH + CO2↑ + MnSO4 + K2SO4 + H2O

CH3-CH2=CH2-CH3 + KMnO4 + H2SO4 → 2CH3-COOH + MnSO4 + H2O